Hello,
I really need help, I tried but I can not get this problem.Practical Application of Quadratic Equations Solve the questions
A rectangular garden has dimensions of 18 feet by 13 feet. A gravel path of uniform width is to be built around the garden. How wide can the path be if there is enough gravel for 516 square feet?
Thank you so much,
Sally


Answer by scott8148(6628)

let x="width of path"

area of garden is 18*13 or 234 __ area of garden plus path is 234+516 or 750

(18+2x)(13+2x)=750 __ 234+62x+4x^2=750 __ dividing by 2 __ 2x^2+31x+117=375 __ subtracting 375 __ 2x^2+31x-258=0

factoring __ (2x+43)(x-6)=0

2x+43=0 __ x=-43/2 __ negative value not realistic

x-6=0 __ x=6